OpenStudy (anonymous):

MEDAL SOLVE (cos Θ − cos Θ)2 + (cos Θ + cos Θ)2

OpenStudy (danjs):

solve what, there is nothing to solve , simplify?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

cos(a)-cos(a) = 0 then expand the second part if those are squared power terms

OpenStudy (anonymous):

so it is 8

OpenStudy (danjs):

\[0^2 + (\cos(\theta)+\cos(\theta))^2=(\cos(\theta)+\cos(\theta))*(\cos(\theta)+\cos(\theta))\]

OpenStudy (danjs):

expand that by doing the foil distribute thing

OpenStudy (anonymous):

4cos^2 theta

OpenStudy (anonymous):

is the answer cos^2 theta

OpenStudy (danjs):

umm

OpenStudy (danjs):

yeah 4 cos squared theata